OpenStudy (anonymous):

physical science : Which of the following would require the greatest number of calories? a. heating 1 g of water from 10°C to 80°C b. heating 10 g of water from 10°C to 40°C c. heating 100 g of water from 10°C to 20°C d. heating 1000 g of water from 10°C to 12°C

OpenStudy (mertsj):

1 calorie is the energy needed to raise the temperature of 1 gram of water through 1 celsius degree. So...a) requires 1(80-10)=70 b) requires 10(40-10)=300 c) requires 100(20-10) = 1000 d) requires 1000(12-10)=2000. So choose d